OpenStudy (anonymous):

An object moves in a straight line so that its velocity after t seconds is v(t). Determine the distance traveled by the object on the given time interval. v(t)=3radical(t) - t on [0,36]...??

OpenStudy (across):

It is a common fact that the integral of the velocity of an object is the distance it travels. Therefore,\[3\int_{0}^{36}\sqrt{t}dt-\int_{0}^{36}tdt.\]Can you integrate that?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

displacement is -216 if that helps...

OpenStudy (across):

Where did you get that answer from?

OpenStudy (anonymous):

someone helped me earlier but didnt help me with distance. and i submitted that answer and it was right..i just dont know how to get the distance

OpenStudy (across):

That's what the above integral evaluates to:\[3\int_{0}^{36}\sqrt{t}dt-\int_{0}^{36}tdt=\left [ 2t^{\frac{3}{2}}\right ]_{0}^{36}-\left [ \frac{t^2}{2}\right ]_{0}^{36}=-216\]

OpenStudy (anonymous):

ohhh so what about distance then?

OpenStudy (across):

Since the distance can't be a negative value, we need to modify the integral as follows:\[\int_{0}^{9}(3\sqrt{t}-t)dt-\int_{9}^{36}(3\sqrt{t}-t)dt=\frac{27}{2}+\frac{459}{2}=243.\]That should be it (if I didn't make a mistake somewhere).

OpenStudy (anonymous):

thank you!

OpenStudy (across):

Was that it?